# On the existence of forward self-similar solutions to the two-dimensional Navier-Stokes equations

### v1

@@ Abstract.
We establish the global existence of forward self-similar solutions to the two-dimensional incompressible Navier-Stokes equations for any divergence-free initial velocity $u_{0}$ that is homogeneous of degree 1-1 and locally Hölder continuous. This result requires no smallness assumption on the initial data. In sharp contrast to the three-dimensional case, where (1)(-1)-homogeneous vector fields are locally square-integrable, the 2D problem is critical in the sense that the initial kinetic energy is locally infinite at the origin, and the initial vorticity fails to be locally integrable. Consequently, the classical local energy estimates are not available. To overcome this, we decompose the solution into a linear part solving the heat equation and a finite-energy perturbation part. By exploiting a kind of inherent cancellation relation between the linear part and the perturbation part, we can control interaction terms and establish the $H^{1}$-estimates for the perturbation part. Further investigating the corresponding Leray system in weighted Sobolev space, we can derive an optimal pointwise estimate. This gives the faster decay of the perturbation part at infinity and enables us to construct global-in-time self-similar solutions.
@@
Navier-Stokes Equation, forward Self-Similar Solution, Large initial data, two-dimensional
@@ 1. Introduction and Main Results
[…]
@@
The study of self-similar solutions is motivated by their fundamental roles in understanding the structure of singularities and the asymptotic behavior of the Navier-Stokes equations. In the context of the regularity problem, (backward) self-similar solutions often serve as natural candidates for blow-up profiles. Although the 2D Navier-Stokes equations are well-known to be globally well-posed for smooth data with finite energy, the forward self-similar solutions with singular initial data provide a critical setting to test the robustness of the theory.
Furthermore, self-similar solutions are quite useful for describing the long-time asymptotics of general solutions, which themselves usually evolve from the singular initial data. For the the Navier-Stokes equations in 2D, it is known that a special class of self-similar solutions called Oseen vortices dominate the long-time behaviour of general solutions to (1.1) with integrable initial vorticity [14, 15, 11]; see also the discussion below. We mention that for the 2D Navier-Stokes equations with integrable vorticity, the solution can have infinite energy in general.
Our main goal is to construct a global self-similar solution for any self-similar initial velocity field, generally with non-integrable vorticity. Before stating our result, we give the definition of the solution we will use; this definition is essentially the same as [26, Definition 8.18].
@@
    $u-e^{t\Delta}u_{0}\in L^{\infty}(0,T;L^{2}(\mathbb{R}^{2}))\cap L^{2,\infty}(0,T;H^{1}(\mathbb{R}^{2}))$ for any $T>0$,
@@
$(e^{t\Delta}u_{0})(x,t)=\int_{\mathbb{R}^{2}}\frac{1}{4\pi t}e^{-\frac{|x-z|^{2}}{4t}}u_{0}(z)dz.$
@@
The main result in this paper is the following theorem.
@@
Assume $u_{0}(x)\in C^{0,\beta}_{loc}$ for some $0<\beta\leq 1$ is a divergence-free, self-similar initial data in $\mathbb{R}^{2}\setminus\{0\}$, and that $\int_{\partial S^{1}}u_{0}\cdot nd\sigma=0$. Suppose that
$A=\|u_{0}\|_{C^{0,\beta}(S^{1})}<+\infty,$
here $S^{1}$ is the unit circle on $\mathbb{R}^{2}$. Then there is a global smooth self-similar solution $u(x,t)\in C^{\infty}(\mathbb{R}^{2}\times(0,\infty))$ to the Cauchy problem (1.1)-(1.2). Moreover, the solution $u$ is an energy perturbed solution in the sense of Definition 1.1. Finally, we have the following pointwise estimates
$|u(x,t)|\leq\frac{C(A,\beta)}{|x|+\sqrt{t}},\quad~|u(x,t)-e^{t\Delta}u_{0}|\leq C(A,\beta)\frac{\sqrt{t}}{|x|^{2}+t},$
@@
Here, $C^{0,\beta}_{loc}$ is the standard local Hölder space and when $\beta=1$, $C^{0,1}_{loc}$ is the local Lipshitz space.
@@
It seems that the condition $\int_{\partial B_{1}}u_{0}\cdot nd\sigma=0$ is necessary for Theorem 1.1 in view of the Jeffery–Hamel solutions [17], which are steady self-similar solutions with generally non-zero flux, and thus are always singular at the origin. We mention that this is different with the 3D case, where any divergence-free ($L^{\infty}_{loc}$) similar vector field $u_{0}$ on $\mathbb{R}^{3}\setminus\{0\}$ automatically satisfies ${\rm div\thinspace}u_{0}=0$ across the origin.
@@
It is possible to reduce the regularity assumptions in Theorem 1.1 to construct large self-similar solutions. See [12, 7, 6, 1] for the results of the 3D case with rough initial data.
@@
We recently got to know that Dallas Albritton, Julien Guillod, Mikhail Korobkov, and Xiao Ren ([24]) obtained a result similar to Theorem 1.1, where the solution at $t=1$ decays as $|x|^{-1-\beta}$ if the initial data $u_{0}$ belongs to $C^{0,\beta}(S^{1})$.
The existence of solutions in a scale-invariant space has been studied extensively, including $L^{n}$, $L^{n,\infty}$, Besov spaces and $BMO^{-1}$ in $n$ dimensions (usually $n=2,3$), typically under smallness assumptions, see [3, 10, 16, 19]. Especially, the small data existence of forward self-similar solutions follows from the uniqueness results in these studies. One may also refer to [26, Chapter 8] and the reference therein for more references.
For $\mathbb{R}^{3}$, recently, the seminal work of Jia and Šverák [18] established the global existence of smooth self-similar solutions for the 3D Navier-Stokes equations with any Hölder continuous self-similar initial data on $\mathbb{R}^{3}\setminus\{0\}$. The main techniques of their approach are to derive the local-in-space Hölder estimates based on the $L_{loc}^{2}$-energy estimates established by Lemarié-Rieusset for the local Leray solutions in [23]. Later in [20], the self-similar solutions were constructed based on the a priori $H^{1}$-estimate obtained for the equivalent Leray equations by Leray’s method of contradiction, which is a kind of blow-up argument. In [4, 5], Bradshaw and Tsai gave a rather direct method to establish the $H^{1}$-estimates for both discretely self-similar solutions and rotated discretely self-similar solutions, and then constructed solutions via the Galerkin approximation. For more related results and constructions of self-similar solutions for the Navier-Stokes equations with more general self-similar initial data, for example, for initial data that lies in $L^{2}_{loc}(\mathbb{R}^{3})$ and Besov space, one can refer to [25, 27, 12, 7, 6, 1]. For the constructions of self-similar solutions for the fractional Navier-Stokes equations and other related models, such as MHD and SQG equation, one can refer to [22, 21, 29, 28, 2].
For $\mathbb{R}^{2}$, note that the (-1)-homogeneous divergence-free velocity field $u_{0}(x)$ has constant circulation on any circle $\partial B_{R}$, hence can also be decomposed as
$u_{0}(x)=\frac{\alpha}{2\pi}\frac{x^{\perp}}{|x|^{2}}+\tilde{u}_{0},$
with $\tilde{u}_{0}$ has zero circulation. Taking curl, we have the corresponding vorticity
$\omega_{0}(x)=\alpha\delta+\textrm{curl }\tilde{u}_{0}.$
When $\tilde{u}_{0}=0$, an explicit class of self-similar solutions taking $\alpha\delta$ as initial data (in vorticity formulation of the Navier-Stokes equations) exists for any $\alpha\in\mathbb{R}$ without smallness restrictions. They are known as Oseen vortices, and the velocity and vorticity are given by:
$u(x,t)=\frac{\alpha}{2\pi}\frac{x^{\perp}}{|x|^{2}}(1-e^{-\frac{|x|^{2}}{4t}}),~\omega(x,t)=\frac{\alpha}{4\pi t}e^{-\frac{|x|^{2}}{4t}}.$
Indeed the vorticity $\omega(x,t)=\frac{\alpha}{4\pi t}e^{-\frac{|x|^{2}}{4t}}$ just solves the heat equation in this case. Oseen vortices characterize the long-time asymptotic behavior of general solutions to (1.1) with integrable initial vorticity [14, 15, 11]. On the other hand, there is no previous known existence result of self-similar solutions for large $\tilde{u}_{0}$, even for the special case $\alpha=0$.
@@
The main difficulty in proving Theorem 1.1, compared to the 3D case, lies in the non-integrability of the initial data, even locally, at the level of both velocity and vorticity on $\mathbb{R}^{2}$. In 2D, a self-similar initial velocity $|u_{0}(x)|\sim|x|^{-1}$ does not belong to $L_{loc}^{2}(\mathbb{R}^{2})$, hence, the local Leray solutions developed in Lemarié-Rieusset in [23] is not applicable in 2D. On the other hand, the initial vorticity does not belong to $L_{loc}^{1}(\mathbb{R}^{2})$ and even can not be viewed as a finite measure as $|\textrm{curl}~u_{0}(x)|\sim|x|^{-2}$. Consequently, we lack a foundational quantity to initiate the estimates at both the velocity and the vorticity level.
@@
To overcome this lack of initial integrability, we adopt a decomposition strategy. We postulate that the singular behavior of the solution is dominated by the linear evolution. Let $e^{t\Delta}u_{0}$ be the caloric lift of the initial data, i.e., the solution to the heat equation starting from the same initial data $u_{0}$. While $e^{t\Delta}u_{0}$ retains the non-$L^{2}$ character of the initial data, it is explicitly computable and smooth for $t>0$. The unknown remainder term $u_{\text{re}}=u-e^{t\Delta}u_{0}$ is then expected to possess better decay properties. By substituting this decomposition into the Navier-Stokes equations and using the self-similarity variables, we derive a perturbed Leray system. The core of our proof lies in establishing global-in-time energy estimates for $u_{\text{re}}$, depending only on initial data. Due to the self-similarity, we can focus on the energy estimates at $t=1$ only.
Our key observation is that there is some inherent cancellation relation between $e^{\Delta}u_{0}$ and $u_{\text{re}}(\cdot,1)$. Using this relation, we derive a crucial identity, see (2.10) below. This identity implies the Dirichlet energy of $u_{\text{re}}(\cdot,1)$ can be controlled by that of the linear part $e^{\Delta}u_{0}$, which is finite by using explicit estimates for $e^{\Delta}u_{0}$. Starting from this Dirichlet energy estimate, we can obtain the $L^{2}$ control for the $u_{\text{re}}(\cdot,1)$ itself via a suitable choice of the multiplier and the interpolation inequality. One can refer to Section 2.2 for a more detailed outline of the proof.
@@
Once we get initial energy estimates for $u_{\text{re}}(\cdot,1)$, we use weighted estimates inspired by [22] to derive pointwise estimate for $u_{\text{re}}(\cdot,1)$. Specifically, we choose appropriate test functions to derive the following key estimate
$\sqrt{1+|x|^{2}}u_{\text{re}}(x,1)\in H^{2}(\mathbb{R}^{2}).$
@@
Based on this regularity and Sobolev embedding, we have the following pointwise estimate of $u_{\text{re}}$ for $x$,
$|u_{\text{re}}(x,1)|\leq\frac{C}{1+|x|}\textrm{ for all }x\in\mathbb{R}^{2}.$
@@
With this decay estimate in hand, we can eventually improve the decay estimates by using the linear theory of Stokes and the heat equation to have that
$|u_{\text{re}}(x,1)|\leq\frac{C}{1+|x|^{2}}\textrm{ for all }x\in\mathbb{R}^{2}.$
@@
This faster decay compared with the linear part $e^{\Delta}u_{0}$, together with standard interior regularity, gives compactness, so that we can employ the Leray-Schauder degree theory to find a self-similar solution to the Navier-Stokes equation (1.1). On the other hand, if $u_{0}\in C^{2}_{loc}(\mathbb{R}^{2})$, we can utilize the computations in [8] to show the optimal decay rate (1.3).
@@
The rest of the paper is organized as follows. We give the proof of the energy estimates for the remainder term $u_{\text{re}}(\cdot,1)=u(\cdot,1)-e^{\Delta}u_{0}$ in Section 2. Next, we derive pointwise decay estimates for $u_{\text{re}}$ by establishing the weighted energy estimates in Section 3. Finally, we prove the main theorem in Section 4 by using the Leray-Schauder degree theory.

### v2

# On the forward self-similar solutions to the two-dimensional Navier-Stokes equations

@@ Abstract.
We establish the global existence of forward self-similar solutions to the two-dimensional incompressible Navier-Stokes equations for any divergence-free initial velocity that is homogeneous of degree 1-1 and locally Hölder continuous. This result requires no smallness assumption on the initial data. In sharp contrast to the three-dimensional case, where (1)(-1)-homogeneous vector fields are locally square-integrable, the major difficulty for the 2D problem is the criticality in the sense that the initial kinetic energy is locally infinite at the origin, and the initial vorticity fails to be locally integrable, so that the classical local energy estimates are not available. Our key ideas are to decompose the solution into a linear part solving the heat equation and a finite-energy perturbation part, and to exploit a kind of inherent cancellation relation between the linear part and the perturbation part. These, together with suitable choices of multipliers, enable us to control the interaction terms and to establish the $H^{1}$-estimates for the perturbation part. Furthermore, we can get an optimal pointwise estimate via investigating the corresponding Leray equations in weighted Sobolev spaces. This gives the faster decay of the perturbation part at infinity and compactness, which play important roles in proving the existence of global-in-time self-similar solutions.
@@
Navier-Stokes equations, Self-similar solutions, Large initial data, two-dimensional
@@ 1. Introduction and Main Results
(1.1) $\left\{\begin{aligned} &\frac{\partial{u}}{\partial t}-\Delta{u}+{u}\cdot\nabla{u}+\nabla\pi=0,\\ &{\rm div\thinspace}{u}=0,\\ \end{aligned}\right.\textrm{ in }\mathbb{R}^{2}\times(0,\infty)$
@@
The study of self-similar solutions is motivated by their fundamental roles in understanding the well-posedness, especially the structure of singularities, the asymptotic behavior, and the uniqueness or non-uniqueness of the Navier-Stokes equations. In the context of the regularity problem, (backward) self-similar solutions often serve as natural candidates for blow-up profiles, and have been excluded in the 3D case ([37, 42]). On the other hand, forward self-similar solutions, which arise from scale-invariant initial data, provide a critical setting to test the robustness of the well-posedness theory. Recently, a lot of important progress has been made on the non-uniqueness of Leray-Hopf solutions through the study of the forward self-similar solutions [26, 3, 21, 23].
The existence of solutions in a scale-invariant space has been studied extensively, including $L^{n}$, $L^{n,\infty}$, Besov spaces and $BMO^{-1}$ in $n$ dimensions (usually $n=2,3$), typically under smallness assumptions, see [7, 14, 20, 28]. Especially, the small data existence of forward self-similar solutions follows from the uniqueness results in these studies. One may also refer to [45, Chapter 8] and the reference therein.
In the three-dimensional case, recently, the seminal work of Jia and Šverák [25] established the global existence of smooth self-similar solutions for the 3D Navier-Stokes equations with any Hölder continuous self-similar initial data on $\mathbb{R}^{3}\setminus\{0\}$. The main techniques of their approach are to derive the local-in-space Hölder estimates based on the $L_{loc}^{2}$-energy estimates established by Lemarié-Rieusset for the local Leray solutions in [35]. Later in [29], the self-similar solutions were constructed based on the a priori $H^{1}$-estimate obtained for the equivalent Leray equations by Leray’s method of contradiction, which is a kind of blow-up argument. In [8, 9], Bradshaw and Tsai gave a rather direct method to establish the $H^{1}$-estimates for both discretely self-similar solutions and rotated discretely self-similar solutions, and then constructed solutions via the Galerkin approximation.
However, the general forward self-similar solutions for two-dimensional Navier-Stokes have neither locally finite energy nor integrable vorticity. This makes the analysis much more challenging. In the two-dimensional case, note that the (-1)-homogeneous divergence-free velocity field $u_{0}(x)$ has constant circulation on any circle, and be decomposed as
(1.3) $u_{0}(x)=\frac{\alpha}{2\pi}\frac{x^{\perp}}{|x|^{2}}+\tilde{u}_{0},$
where $\tilde{u}_{0}$ has zero circulation. In the case $\tilde{u}_{0}=0$, an explicit class of self-similar solutions is known as Oseen vortices, whose velocity and vorticity are given by
$u(x,t)=\frac{\alpha}{2\pi}\frac{x^{\perp}}{|x|^{2}}(1-e^{-\frac{|x|^{2}}{4t}})\,\,\text{ and }\,\,~\omega(x,t)=\frac{\alpha}{4\pi t}e^{-\frac{|x|^{2}}{4t}},$
respectively. In fact, the Oseen vortices characterize the long-time asymptotic behavior of general solutions to (1.1) with integrable initial vorticity [18, 19, 15]. Therefore, the forward self-similar solutions also play crucial roles in understanding the long-time behavior of general solutions of (1.1). However, there is no known existence result of self-similar solutions with the intial data of the form (1.3) with large $\tilde{u}_{0}$, even for the special case $\alpha=0$.
1.1. Main results
Our main goal is to construct a global forward self-similar solution for any self-similar initial velocity field, generally with non-integrable vorticity. Before stating our result, we give the definition of the solution to (1.1), which is essentially the same as [45, Definition 8.18].
@@ 1.1. Main results
    $u-e^{t\Delta}u_{0}\in L^{\infty}(0,T;L^{2}(\mathbb{R}^{2}))\cap L^{2}(0,T;H^{1}(\mathbb{R}^{2}))$ for any $T>0$,
@@
$(e^{t\Delta}u_{0})(x)=\int_{\mathbb{R}^{2}}\frac{1}{4\pi t}e^{-\frac{|x-z|^{2}}{4t}}u_{0}(z)dz.$
@@
The main results in this paper are as follows.
@@
Assume $u_{0}(x)\in C^{0,\beta}_{loc}$ for some $0<\beta\leq 1$ is a divergence-free, self-similar initial data in $\mathbb{R}^{2}\setminus\{0\}$, and that $\int_{S^{1}}u_{0}\cdot nd\sigma=0$. Here, $S^{1}$ is the unit circle on $\mathbb{R}^{2}$ and $C^{0,\beta}_{loc}$ is the standard local Hölder space, when $\beta=1$, $C^{0,1}_{loc}$ is the local Lipshitz space. Suppose that
$A=\|u_{0}\|_{C^{0,\beta}(S^{1})}<+\infty.$
Then the following statements hold.
  • (i)
    There exists at least one global smooth self-similar solution $u(x,t)\in C^{\infty}(\mathbb{R}^{2}\times(0,\infty))$ to the Cauchy problem (1.1)-(1.2), which is indeed an energy perturbed solution in the sense of Definition 1.1.
  • (ii)
    The following pointwise estimates hold
    (1.4) $|u(x,t)|\leq\frac{C(A,\beta)}{|x|+\sqrt{t}},\quad~|u(x,t)-e^{t\Delta}u_{0}|\leq C(A,\beta)\frac{\sqrt{t}}{|x|^{2}+t},$
@@
There are a few remarks in order.
Remark 1.1.
It is interesting to note that any divergence-free self-similar vector field $u_{0}\in L^{\infty}_{loc}(\mathbb{R}^{3}\setminus\{0\})$ automatically satisfies ${\rm div\thinspace}u_{0}=0$ across the origin. However, this property does not hold for a two-dimensional vector field. For example,
$u_{0}(x)=\frac{1}{2\pi}\frac{x}{|x|^{2}}$
is divergence-free on $\mathbb{R}^{2}\setminus\{0\}$ but ${\rm div\thinspace}u_{0}=\delta$ in the sense of distributions on $\mathbb{R}^{2}$. Since the Navier-Stokes equations ask the velocity field to be divergence-free for $t>0$, it is natural to require the initial data $u_{0}$ to be divergence-free in $\mathbb{R}^{2}$ at least in the sense of distributions, i.e., $\int_{S^{1}}u_{0}\cdot nd\sigma=0$.
Remark 1.2.
Together with the techniques developed in [44, 47, 16, 11, 10, 1], where large self-similar solutions were constructed with initial data that lie in $L^{2}_{loc}(\mathbb{R}^{3})$ and Besov space, etc, we may reduce the regularity assumptions in Theorem 1.1 to construct two-dimensional forward self-similar solutions with general initial data.
Remark 1.3.
The leading term $e^{t\Delta}u_{0}$ satisfies
$\left|e^{t\Delta}u_{0}(x)\right|\leq\frac{C(A)}{|x|+\sqrt{t}}.$
The details are referred to Lemma 2.1.
@@
In recent years, the self-similar solutions of the steady Navier-Stokes equations were also studied extensively, both on the rigidity part and the existence part. For example, nontrivial self-similar solutions were obtained in [22, 24, 34]: Hamel solutions when the dimension $n=2$ and Landau solutions when $n=3$. The rigidity of Hamel solutions and Landau solutions in the class of self-similar solutions was established in [41, 40]. The importance of these self-similar solutions in understanding singularity and far-field behavior of solutions to the steady Navier-Stokes equations was illustrated in [36, 30] and references therein. There is also a great interest in studying the high-dimensional steady Navier-Stokes equations, and it is known that for dimensions $n\geq 4$, all self-similar solutions are trivial (see [43], [40, Corollary in Section 2]). This rigidity result has been extended to all solutions with scale-invariant bounds $C/|x|$ without any smallness assumption on $C$ in [5]. The existence of self-similar solutions to steady Navier-Stokes equations with small scale-invariant forces has been obtained in [31, 38, 27, 46] for $n\geq 3$. Recently, the existence of self-similar solutions with arbitrarily large scale-invariant forces has been proved in [6] up to dimension 16.
@@
For the constructions of self-similar solutions for the fractional Navier-Stokes equations and other related models, such as MHD and SQG equation, one can refer to [33, 32, 49, 48, 2]. It is also possible to study forward self-similar solutions for the associated 2D system with the aid of the methods developed in this paper.
@@
We recently got to know that Dallas Albritton, Julien Guillod, Mikhail Korobkov, and Xiao Ren ([4]) obtained a result similar to Theorem 1.1, where $u(x,t)-e^{t\Delta}u_{0}$ at $t=1$ admits spatial decay $|x|^{-1-\beta}$ if the initial data $u_{0}$ belongs to $C^{0,\beta}(S^{1})$. After both papers appeared on ArXiv, we noticed that the methods are quite different. We use straightforward energy estimates and do not use special quantities like the Bernoulli pressure.
1.2. Main difficulties and the key ideas
@@
The main difficulty in proving Theorem 1.1, compared to the 3D case, lies in the non-integrability of the initial data, even locally, at the level of both velocity and vorticity in $\mathbb{R}^{2}$. In 2D, a self-similar initial velocity $|u_{0}(x)|\sim|x|^{-1}$ does not belong to $L_{loc}^{2}(\mathbb{R}^{2})$, hence, the local Leray solutions developed in [35] are not applicable. On the other hand, the initial vorticity does not belong to $L_{loc}^{1}(\mathbb{R}^{2})$ and even cannot be viewed as a finite measure as $|\textrm{curl}~u_{0}(x)|\sim|x|^{-2}$. Consequently, we lack a fundamental quantity to initiate the estimates at both the velocity and the vorticity level.
@@
To overcome this lack of initial integrability, we adopt a decomposition strategy. We postulate that the singular behavior of the solution is dominated by the linear evolution. Let $e^{t\Delta}u_{0}$ be the caloric lift of the initial data, i.e., the solution to the heat equation starting from the same initial data $u_{0}$. While $e^{t\Delta}u_{0}$ retains the non-$L^{2}$ character of the initial data, it is explicitly computable and smooth for $t>0$. The unknown remainder term $u_{\text{re}}=u-e^{t\Delta}u_{0}$ is then expected to possess better decay properties. The core of our proof lies in establishing global-in-time energy estimates for $u_{\text{re}}$, depending only on initial data.
Due to the self-similarity, we only need to focus on the energy estimates at $t=1$. To do so, we use the self-similar variables and substitute this decomposition into the Navier-Stokes equations, thereby deriving the perturbed Leray equations. More precisely, we let
$u(x,t)=\frac{1}{\sqrt{t}}v\left(\frac{x}{\sqrt{t}}\right)=\frac{1}{\sqrt{t}}v(y).$
and
$v_{0}(y)=e^{\Delta}u_{0}(y),~v_{\text{re}}=v-v_{0}.$
We can derive that (see Section 2.1)
$\left\{\begin{aligned} &-\Delta v_{\text{re}}-\frac{1}{2}v_{\text{re}}-\frac{1}{2}y\cdot\nabla v_{\text{re}}+v_{0}\cdot\nabla v_{\text{re}}+v_{\text{re}}\cdot\nabla v_{0}+v_{\text{re}}\cdot\nabla v_{\text{re}}+\nabla q=-v_{0}\cdot\nabla v_{0},\\ &{\rm div\thinspace}v_{\text{re}}=0,\end{aligned}\right.\textrm{ in }\mathbb{R}^{2}.$
The energy estimates of $v_{\text{re}}$ cannot be achieved directly if one tests the equation of $v_{\textrm{re}}$ above by $v_{\text{re}}$. This is a key difference compared with the three-dimensional case, where there is an additional term $\frac{1}{4}\int_{\mathbb{R}^{2}}|v_{\text{re}}|^{2}$ helps to close the energy estimates; see [8, 9]. The difficulty at this stage is how to derive some initial, although maybe insufficient, estimates for $v_{\text{re}}$.
Our first key observation is that one can control the Dirichlet energy $\|\nabla v_{\text{re}}\|_{L^{2}(\mathbb{R}^{2})}^{2}$ of $v_{\text{re}}$ by using some inherent cancellation relation between $v_{0}$ and $v_{\text{re}}$. Indeed, this relation leads to a crucial identity (see also (2.13) below), specific to 2D, which is
$\int_{\mathbb{R}^{2}}|\nabla v_{\text{re}}|^{2}dy+2\int_{\mathbb{R}^{2}}\nabla v_{0}\cdot\nabla v_{\text{re}}dy=0.$
And one sees that $\|\nabla v_{\text{re}}\|_{L^{2}(\mathbb{R}^{2})}$ can be controlled by $\|\nabla v_{0}\|_{L^{2}(\mathbb{R}^{2})}$, which is finite by using explicit estimates for $e^{\Delta}u_{0}$. This Dirichlet energy estimate $\|\nabla v_{\text{re}}\|_{L^{2}(\mathbb{R}^{2})}$ for $v_{\text{re}}$ is the starting point of our subsequent analysis.
Our next important observation is that the convection term $\frac{1}{2}y\cdot\nabla v_{\text{re}}$ can be more effectively used if we choose the multiplier $|v_{\text{re}}|^{p-2}v_{\text{re}}$, where $1<p<2$. Testing the equation of $v_{\textrm{re}}$ above, the linear term $\frac{1}{2}v_{\text{re}}-\frac{1}{2}y\cdot\nabla v_{\text{re}}$ gives a positive quantity
$\int_{\mathbb{R}^{2}}(-\frac{1}{2}v_{\text{re}}-\frac{1}{2}y\cdot\nabla v_{\text{re}})\cdot|v_{\text{re}}|^{p-2}v_{\text{re}}=\int_{\mathbb{R}^{2}}\left(\frac{1}{p}-\frac{1}{2}\right)|v_{\text{re}}|^{p}dy.$
This quantity, together with the estimate of $\|\nabla v_{\text{re}}\|_{L^{2}(\mathbb{R}^{2})}$ is essential for us to close the $L^{p}$-estimate of $v_{\text{re}}$ for any $1<p<2$. This also implies that $v_{\textrm{re}}$ decays like $\frac{1}{1+|y|^{2}}$ in an average sense. The application of interpolation inequalities leads to the $L^{2}$-estimate, and hence, $H^{1}$-estimate of $v_{\text{re}}$. One can refer to Section 2.2 for a more detailed outline of the proof.
@@
Once we get the $H^{1}$-estimate for $v_{\text{re}}$, we proceed to establish the weighted energy estimates and to derive pointwise estimates for $v_{\text{re}}$ inspired by the methods developed in [33]. Specifically, we choose the test function $|y|^{2}v_{\text{re}}$, and the linear term $\frac{1}{2}v_{\text{re}}-\frac{1}{2}y\cdot\nabla v_{\text{re}}$ contributes a crucial positive term
$\int_{\mathbb{R}^{2}}\left(-\frac{1}{2}v_{\text{re}}-\frac{1}{2}y\cdot\nabla v_{\text{re}}\right)\cdot(|y|^{2}v_{\text{re}})\,dy=\frac{1}{2}\int_{\mathbb{R}^{2}}|y|^{2}|v_{\text{re}}|^{2}\,dy.$
@@
This integral, due to the convection $\frac{1}{2}y\cdot\nabla v_{\text{re}}$, together with $H^{1}$-estimate of $v_{\text{re}}$ leads to $H^{1}$-estimate of $|y|v_{\text{re}}$. After bootstrapping to $H^{2}$-estimate of $|y|v_{\text{re}}$, we have the following pointwise estimate of $v_{\text{re}}$ by Sobolev embedding,
$|v_{\text{re}}(y)|\leq\frac{C}{1+|y|}\textrm{ for all }y\in\mathbb{R}^{2}.$
@@
With this decay estimate in hand, we can further improve the decay estimates by using the linear theory of Stokes and the heat equation to have that
$|v_{\text{re}}(y)|\leq\frac{C}{1+|y|^{2}}\textrm{ for all }y\in\mathbb{R}^{2}.$
@@
This faster decay compared with the linear part $e^{\Delta}u_{0}$, together with standard interior regularity, gives compactness, so that we can employ the Leray-Schauder degree theory to find a solution $v_{\textrm{re}}$, hence, a self-similar solution to the Navier-Stokes equation (1.1). On the other hand, if $u_{0}\in C^{2}_{loc}(\mathbb{R}^{2})$, we can utilize the computations in [12] to show the optimal decay rate (1.5).
@@
The rest of the paper is organized as follows. We prove the energy estimates for the remainder term $u_{\text{re}}(\cdot,1)=u(\cdot,1)-e^{\Delta}u_{0}$ in Section 2. Next, we establish the weighted energy estimates and derive pointwise decay estimates for $u_{\text{re}}(\cdot,1)$ in Section 3. Finally, the main theorem is proved in Section 4 by using the Leray-Schauder degree theory.